# ChartAttack: Testing the Vulnerability of LLMs to Malicious Prompting in Chart Generation

### v1

@@ Abstract
Multimodal large language models (MLLMs) are increasingly used to automate chart generation from data tables, enabling efficient data analysis and reporting but also introducing new misuse risks. In this work, we introduce ChartAttack, a novel framework for evaluating how MLLMs can be misused to generate misleading charts at scale. ChartAttack injects misleaders into chart designs, aiming to induce incorrect interpretations of the underlying data. Furthermore, we create AttackViz, a chart question-answering (QA) dataset where each (chart specification, QA) pair is labeled with effective misleaders and their induced incorrect answers. Experiments in in-domain and cross-domain settings show that ChartAttack significantly degrades the QA performance of MLLM readers, reducing accuracy by an average of 19.6 points and 14.9 points, respectively. A human study further shows an average 20.2 point drop in accuracy for participants exposed to misleading charts generated by ChartAttack. Our findings highlight an urgent need for robustness and security considerations in the design, evaluation, and deployment of MLLM-based chart generation systems. We make our code and data publicly available 111[Uncaptioned image]https://github.com/insait-institute/chartAttack.
@@ 1 Introduction
Charts are widely used to communicate complex information across various domains, including political, social, environmental, and health, (Lauer and O’Brien, 2020; Huang et al., 2025). They play a critical role during crises, such as the COVID-19 pandemic (Zhang et al., 2021; Woloshin et al., 2023). However, poorly designed or intentionally manipulated charts can propagate misinformation (Huff and Geis, 1993; Lan and Liu, 2025). Misleading charts distort the interpretation of the underlying data through misleading techniques, also known as misleaders, which are design choices that violate established visualization principles in ways that systematically bias perception or inference, such as inverting axes to reverse perceived trends. Prior work has shown that misleading charts can significantly decrease the performance of both human readers (Pandey et al., 2014, 2015; O’Brien and Lauer, 2018; Yang et al., 2021; Ge et al., 2023; Rho et al., 2024) and MLLMs (Bharti et al., 2024; Bendeck and Stasko, 2025; Chen et al., 2025; Tonglet et al., 2025a) in a QA setting.
@@
Chart creation has been democratized via user-friendly tools and social media (Pandey et al., 2015), and designers increasingly use MLLMs for chart generation and analysis Shen et al. (2024); Ahn and Kim (2025). While MLLMs simplify legitimate tasks, they can be exploited to generate misleading content at scale (Pan et al., 2023; Sallami et al., 2024; Zugecova et al., 2025), including misleading charts (Figure 1). However, the effectiveness of MLLM-based misleading chart generation and its impact on readers have not been systematically quantified.
@@
In this work, we present the first systematic study of this jailbreaking attack (Wei et al., 2023; Lin et al., 2024). We introduce ChartAttack (Figure 2), a framework that automatically applies misleaders to chart annotations with the objective of deceiving readers with respect to a specific question about the chart. Chart annotations are JSON files containing the data table and basic formatting specifications needed to generate the chart. ChartAttack applies known misleaders that alter a chart’s design without changing the underlying data, allowing deliberate generation of misleading charts that remain data-consistent. To support this task, we introduce AttackViz, a multi-label chart QA dataset covering horizontal and vertical bar charts as well as line charts. Each instance contains chart annotations, an associated question, and a set of misleaders with annotations specifying how each is applied and the incorrect answers it causes.
@@
We evaluate ChartAttack on both MLLM and human readers. Our model reduces average MLLM QA accuracy by 19 percentage points (pp) on AttackViz. Our experiments also show that ChartAttack generalizes to other datasets, reducing accuracy by an average of 14.9 pp in cross-dataset evaluation, and by 13.3 pp for human participants, demonstrating the effectiveness of misleading visualization attacks both on MLLMs and humans.
@@
We summarize our contributions as follows: (1) We introduce ChartAttack, the first model for automatically generating misleading charts via systematically applied misleaders, which can be precisely specified, reproduced, and parameterized to induce specific misinterpretations in charts. (2) We present AttackViz, a chart QA dataset with structured annotations that include both the original chart annotations and correct answers, as well as the modified chart annotations with applied misleaders and the resulting incorrect answers. (3) We provide an extensive evaluation of misleading visualization attacks on both MLLMs and human readers.
@@ 3 ChartAttack framework
ChartAttack (Figure 2) is a framework that generates misleading charts by applying misleaders to chart annotations, with the goal of inducing incorrect responses to questions associated with the chart. The input consists of chart annotations with data and basic formatting specifications, along with a question and its correct answer. The framework has two components. The Demonstration Selection module retrieves similar examples and uses them as demonstrations in a few-shot prompting setup. The Misleading Generator module takes the chart annotations, the question and correct answer, and the retrieved demonstrations. It outputs a list of items, each specifying a selected misleader, a modified annotation snippet applying the technique, and a misleading answer. Each misleading answer is plausible but incorrect and uses the same type and units as the correct answer.
3.1 Demonstration selection module
@@
The effectiveness of in-context learning strongly depends on the quality of selected examples (Liu et al., 2022; Wang et al., 2024). To retrieve relevant demonstrations from a large corpus, we fine-tune an SBERT model Reimers and Gurevych (2019) using Multiple Negative Ranking Loss Henderson et al. (2017). A demonstration–input pair is considered positive if their sets of misleaders match exactly. For both corpus candidates and input instances, SBERT encodes the concatenation of the question and its chart JSON annotation. Top-$k$ demonstrations are retrieved using cosine similarity and included in the prompt of the Misleading Generator module to guide misleading chart generation. We train a separate retriever for each chart type because each type is affected by a different set of misleaders and has distinct chart semantics. Experimental results are reported in § 6.1, and the dataset creation process for this module is detailed in Appendix A.
3.2 Misleader-generator module
@@
The module applies misleaders to chart JSON annotations to induce incorrect answers to associated questions. We use code-based instruction-tuned MLLMs to modify chart JSON annotations, as they outperform general MLLMs on structured reasoning tasks (Madaan et al., 2022). We select models based on their performance on the Human-Eval benchmark and use a few-shot prompting strategy with five demonstrations. Oracle results in Appendix B indicate strong gains up to 3-shot prompting, with 5-shot yielding the best overall Micro and Macro F1-scores. The module takes three inputs: (1) chart annotations, containing the data and basic formatting specifications, (2) the associated question, and (3) similar examples retrieved by the Demonstration Selection module.
@@
We use a prompt template for all chart types. In a single inference step, the model follows a structured multi-step reasoning process: (i) select misleaders compatible with the chart and context; (ii) specify minimal modifications to apply each misleader without altering other elements; and (iii) produce a misleading answer based on the applied misleader. This design ensures consistent generation of misleading chart variants. Details are provided in Appendix D.
4 The AttackViz Corpus
@@ 4.2 Data Preprocessing
4.3 Rule-based misleader chart generation
We generate misleading charts using a rule-based system. The system implements eight misleaders selected from the 74 categories in Lo et al. (2022). Table 1 lists the definitions of the selected techniques. We select misleaders based on two criteria: their frequency in real-world examples (Lo et al., 2022) and prior studies on QA with misleading charts (Ge et al., 2023; Bharti et al., 2024) and designer support (Lo et al., 2023). We implement the system in Python using Matplotlib (Hunter, 2007). The system modifies the chart JSON annotations to apply the selected misleader, then parses the annotations to generate the corresponding chart image. Using the same process without applying a misleader produces the correct chart. Operating at the annotation level allows the pipeline to adapt to other visualization libraries.
@@ 4.3 Rule-based misleader chart generation
4.4 Evaluation and filtering process
We perform chart QA using each correct chart and its misleading counterpart from our rule-based system. We measure performance with relaxed accuracy Masry et al. (2022); Methani et al. (2020). We select three instruction-tuned MLLMs based on their ChartQA test set performance (Masry et al., 2022): QwenVL 2.5-32B Bai et al. (2025), InternVL 3.0-38B Zhu et al. (2025), and KimiVL-A3B Team et al. (2025). We filter the dataset to ensure high-quality instances. We retain cases where most models answer correctly on the correct chart and incorrectly on the misleading chart. We then apply a consistency filter to confirm that errors result from the misleader: numeric answers must have a standard deviation below 0.5, and textual answers must have a majority identical incorrect response. We determine the final misleading answer by averaging incorrect numeric responses or taking the majority vote for textual responses.
4.5 Cross-domain extension
We apply the same pipeline to ChartQA (Masry et al., 2022), a dataset of charts from real-world sources such as Statista, Pew Research Center, Our World in Data, and the OECD. Due to inconsistent or missing annotations, we merge all instances into a single test set. We provide a detailed explanation of these inconsistencies and the original size of ChartQA in Appendix C.1.
@@ 5 Experiments
Figure 4: Average accuracy on AttackViz. Top: Results by model. Bottom: Results by misleader. Colors indicate setting and dataset: PlotQA: Accuracy on correct charts and Accuracy on misleading charts. ChartQA: Accuracy on correct charts and Accuracy on misleading charts
@@ 5.1 Experimental setup
We perform all experiments on the test splits of AttackViz, derived from PlotQA and ChartQA. We evaluate both in-domain and cross-domain generalization. In the in-domain setting, demonstrations and test instances come from the same source dataset (PlotQA). In the cross-domain setting, demonstrations are selected from the PlotQA train split while test instances come from ChartQA. For each test instance, the Demonstration Selection module retrieves the most relevant training examples to use as demonstrations.
@@
Following prior work on evaluating MLLMs vulnerabilities to misleading charts Tonglet et al. (2025a), we evaluate 11 open-weight, instruction-tuned models from three families. The models include Ovis-2.5 (2B, 9B) Lu et al. (2025), InternVL-3.5 (1B, 2B, 4B, 8B, 14B, 38B) Wang et al. (2025), and LLaVA-1.6 (7B, 13B, 34B) Liu et al. (2024). We use the HuggingFace Transformers library Wolf et al. (2019) to load the models and run inference.
@@
We evaluate each model under two configurations: (i) with the correct chart and (ii) with the misleading chart generated by ChartAttack. Following prior work Methani et al. (2020); Masry et al. (2022), we report relaxed accuracy as the primary metric. We also define two deception-rate metrics. Deception rate (originally correct) captures the fraction of answers that were correct on the correct chart but that the model switches to the misleading answer. Deception rate (originally incorrect) captures the fraction of answers that were incorrect on the correct chart but that the model produces the misleading answer, showing whether misleading charts reinforce existing errors. High deception rates are challenging, as these metrics count only exact switches, requiring the Misleader-Generator to precisely anticipate the victim’s behavior.
@@ 5.2 MLLM-based evaluation results
All models perform worse on misleading charts compared to correct charts, with drops ranging from 4.9 to 37.5 pp and an average drop of 19.7 pp. Lower-accuracy models experience smaller decreases (5–9 pp), whereas higher-performing models show larger drops (23–38 pp). Conditional deception rates indicate that these drops primarily result from originally correct answers being shifted to attacker-generated misleading answers (average 11.1%), while originally incorrect answers are largely unaffected (average 2.1%). The impact varies across model sizes and families: LLaVa 1.6 models drop from 4.9 pp (7B) to 9.2 pp (34B), showing only a modest increase despite nearly a five-fold size difference, whereas Intern VL-3.5 models range from 15.6 pp (1B) to 37.5 pp (8B) and 25.4 pp (38B), suggesting that vulnerability does not scale monotonically with model size. Comparisons of similar-size models also reveal variability: InternVL-3.5 14B (26.1 pp) and Ovis 2.5 9B (24.4 pp) exhibit comparable drops. These patterns suggest that factors beyond size influence susceptibility to misleading charts.
Figure 5: Average deception rate on AttackViz. Top: Results by model. Bottom: Results by misleader. Colors indicate setting and dataset: PlotQA: Deception rate (DR) on correct and Deception rate (DR) on incorrect. ChartQA: Deception rate (DR) on correct and Deception rate (DR) on incorrect.
At the misleader level, techniques that substantially alter visual perception, such as Inappropriate use of stacked, 3D, and Inappropriate use of log scale, are associated with the largest reductions in accuracy on misleading charts, reaching 26.4%, 32.7%, and 43.2%, respectively. These correspond to performance drops of 40.2 pp, 32.6 pp, and 27.0 pp, as well as higher deception rates on originally correct answers (20.1%, 11.3%, and 7.2%). Misrepresentation, Inverted axis, and Inappropriate use of line produce moderate drops of 19.7-15 pp and deception rates of 10.8-8.1%. In contrast, Ineffective color scheme has minimal effect, with no accuracy drop and a 0.5% deception rate. The dual axis technique causes only a small overall performance drop (2.2 pp), but in the few cases where it is applied effectively, it can still generate the misleading answers produced by ChartAttack, with a notable deception rate of 10.3%.
@@
Consistent with results from in-domain experiments, all models achieve lower accuracy on misleading charts compared to correct charts, with decreases ranging from 5.1 to 19.6 pp, resulting in an average drop of 14.9 pp. Stronger models are more affected: InternVL-3.5 38B decreases by 18.3 pp, Ovis-2.5 9B by 14 pp, and InternVL-3.5 14B by 18.6 pp. Conditional deception rates remain low across all models (average 3.3% for originally correct answers, 1.3% for originally incorrect answers), indicating that attacker-generated misleading answers rarely convert originally correct predictions in cross-domain settings.
At the technique level, several trends observed in the in-domain experiments persist. 3D remains the most impactful one, reducing accuracy to 25.3% (a 58.3 pp drop) with a deception rate on originally correct answers of 4.6%. Misrepresentation is the second-most effective one, with 55.5% accuracy, a 21.8 pp decrease, and a 6.2% deception rate on originally correct answers. Inappropriate use of stacked bars also remains effective, with 57.6% accuracy, an 18.9 pp decrease, and a 4% deception rate on originally correct answers. In contrast, Dual axis and Ineffective color scheme remain largely unsuccessful. Dual axis produces a negligible accuracy increase of 0.1 pp, while ineffective color scheme cause a small 2.2 pp drop. Their impact is primarily on answers that are already incorrect with the correct chart, with deception rates on originally incorrect answers of 0.7% and 2.0%, respectively. Inappropriate use of lines and log scales achieve modest performance drops of 0.8 pp and 4.9 pp, respectively. These results suggest that while some misleaders generalize across domains, others are more sensitive to chart and question semantics.

### v2

@@ Abstract
Multimodal large language models (MLLMs) are increasingly used to automate chart generation from data tables, enabling efficient data analysis and reporting but also introducing new misuse risks. In this work, we introduce ChartAttack, a novel framework for evaluating how MLLMs can be misused to generate misleading charts at scale. ChartAttack injects misleaders into chart designs, aiming to induce incorrect interpretations of the underlying data. Furthermore, we create AttackViz, a chart question-answering (QA) dataset where each (chart specification, QA) pair is labeled with effective misleaders and their induced incorrect answers. ChartAttack significantly degrades QA performance, reducing MLLM accuracy by 17.2 points in-domain and 11.9 cross-domain. Preliminary human results (limited sample size) indicate a 20.2-point accuracy drop. Finally, we demonstrate that AttackViz can be used to fine-tune MLLMs to improve robustness against misleading charts. Our findings highlight an urgent need for robustness and security considerations in the design, evaluation, and deployment of MLLM-based chart generation systems. We make our code and data publicly available 111[Uncaptioned image] https://github.com/insait-institute/chartAttack.
@@ 1 Introduction
Charts are widely used to communicate complex information across various domains, including political, environmental, and health domains (Lauer and O’Brien, 2020; Huang et al., 2025). They play a critical role during crises, such as the COVID-19 pandemic (Zhang et al., 2021; Woloshin et al., 2023). However, poorly designed or intentionally manipulated charts can propagate misinformation (Huff and Geis, 1993; Lan and Liu, 2025). Misleading charts distort the interpretation of the underlying data through misleading techniques, also known as misleaders, which are design choices that violate established visualization principles in ways that systematically bias perception or inference, such as inverting axes to reverse perceived trends. Prior work has shown that misleading charts can significantly decrease the performance of both human readers (Pandey et al., 2014, 2015; O’Brien and Lauer, 2018; Yang et al., 2021; Ge et al., 2023; Rho et al., 2024) and MLLMs (Bharti et al., 2024; Bendeck and Stasko, 2025; Chen et al., 2025; Tonglet et al., 2025a) in a QA setting.
@@
Chart creation has been democratized via user-friendly tools (Pandey et al., 2015), and designers increasingly use MLLMs for chart generation and analysis Shen et al. (2024); Ahn and Kim (2025). While MLLMs simplify legitimate tasks, they can be exploited to generate misleading content at scale (Pan et al., 2023; Sallami et al., 2024; Zugecova et al., 2025), including misleading charts (Figure 1). However, the effectiveness of MLLM-based misleading chart generation and its impact on readers have not been systematically quantified.
@@
In this work, we present the first systematic study of this jailbreaking attack (Wei et al., 2023; Lin et al., 2024). We introduce ChartAttack (Figure 2), a framework that automatically applies misleaders to chart annotations with the objective of deceiving readers with respect to a specific question about the chart. Chart annotations are JSON files containing the data table and basic formatting specifications needed to generate the chart. ChartAttack applies known misleaders that alter a chart’s design without changing the underlying data, allowing deliberate generation of misleading charts that remain data-consistent. To support this task, we introduce AttackViz, a multi-label chart QA dataset covering bar (horizontal and vertical) and line charts. Each instance contains chart annotations, an associated question, and a set of misleaders with annotations specifying how each is applied and the incorrect answers it causes.
@@
We evaluate ChartAttack on both MLLMs and human readers. It reduces average MLLM QA accuracy by 17.2 percentage points (pp) in-domain and 11.9 pp cross-domain. In a preliminary human study, misleading charts generated by ChartAttack reduce participant accuracy by 13.3 pp, demonstrating the effectiveness of misleading visualization attacks for both MLLMs and humans. We further show that AttackViz can be used to fine-tune an MLLM for improved robustness to misleading charts, increasing test-set performance by 8.4 pp.
@@
We summarize our contributions as follows: (1) We introduce ChartAttack, the first framework for automatically generating misleading charts through systematically applied misleaders that can be precisely specified, reproduced, and parameterized to induce targeted misinterpretations. (2) We present AttackViz, a chart QA dataset with structured annotations containing both original chart annotations with correct answers and modified annotations with applied misleaders and the resulting incorrect answers. (3) We provide an extensive evaluation of misleading visualization attacks on both MLLMs and human readers and demonstrate the potential of fine-tuning MLLMs to improve robustness.
@@ 3 ChartAttack framework
ChartAttack (Figure 2) is a framework for generating misleading charts by applying misleaders to chart annotations to induce incorrect answers to chart-related questions. The input consists of chart annotations with data and basic formatting specifications, along with a question and its correct answer. The framework has two components. The Demonstration selection module retrieves similar examples and uses them as demonstrations in a few-shot prompting setup. The Misleading Generator module takes the chart annotations, the question and correct answer, and the retrieved demonstrations. It outputs a list of items, each specifying a selected misleader, a modified annotation snippet applying the technique, and a misleading answer. Each misleading answer is plausible but incorrect and uses the same type and units as the correct answer.
Demonstration selection module.
@@
The effectiveness of in-context learning depends on the quality of selected examples (Liu et al., 2022; Wang et al., 2024). To retrieve relevant demonstrations from a large corpus, we fine-tune an SBERT model Reimers and Gurevych (2019) using Multiple Negative Ranking Loss Henderson et al. (2017). A demonstration–input pair is considered positive if their sets of misleaders match exactly. For both corpus candidates and input instances, SBERT encodes the concatenation of the question and its chart JSON annotation. Top-$k$ demonstrations are retrieved using cosine similarity and included in the prompt of the Misleading Generator module to guide misleading chart generation. We train a separate retriever for each chart type because each type is affected by a different set of misleaders and has distinct chart semantics. Experimental results are reported in § 4, and the dataset creation process for this module is detailed in Appendix A.
Misleader-generator module.
@@
The module applies misleaders to chart JSON annotations to induce incorrect answers to associated questions. We use code-based instruction-tuned MLLMs to modify chart JSON annotations, as they outperform general MLLMs on structured reasoning tasks (Madaan et al., 2022). We select models based on their performance on the Human-Eval benchmark Zhang et al. (2025) and use a few-shot prompting strategy with $k$ demonstrations. The module takes three inputs: (1) chart annotations, containing the data and basic formatting specifications, (2) the associated question, and (3) similar examples retrieved by the Demonstration Selection module. We conduct ablation studies on the demonstration selection and misleader-generator modules to identify the optimal loss function, downsampling strategy, code-based instruction MLLM, retrieval strategy, and number of few-shot demonstrations. Detailed results are reported in Appendix B.
@@
We use a prompt template for all chart types. In a single inference step, the model follows a structured multi-step reasoning process: (i) select misleaders compatible with the chart and context; (ii) specify minimal modifications to apply each misleader without altering other chart elements; and (iii) produce a misleading answer based on the applied misleader. This design ensures consistent generation of misleading chart variants. Details are provided in Appendix D.
4 AttackViz corpus
@@
Rule-based misleading chart generation and chart coverage.
We generate misleading charts using a rule-based system implementing 11 misleaders from the taxonomy of Lo et al. (2022). Table 1 summarizes the selected techniques. Misleaders are chosen according to six criteria: (1) at least five occurrences in real-world examples; (2) previously studied in misleading chart QA (Ge et al., 2023; Bharti et al., 2024) or visualization design-support research (Lo et al., 2023); (3) the correct answer to the associated question remains unchanged after applying the misleader; (4) the technique violates grammar rules; (5) the underlying data table remains correct; and (6) the misleader can be implemented in Python. A detailed table indicating which criteria each taxonomy misleader satisfies is provided in Appendix C.1. We focus on bar (horizontal and vertical) and line charts, which dominate the taxonomy (64%) and account for 49% of real-world misleading visualizations in the MisViz benchmark Tonglet et al. (2025b). While ChartQA and ChartX include additional chart types, these are affected by fewer misleaders. For example, pie charts are affected by two and map charts by three of the eleven techniques. Consequently, these chart types are less relevant due to their smaller misleading-design search space. We implement the system in Python using Matplotlib (Hunter, 2007). The system modifies chart JSON annotations to apply a misleader and then parses the annotations to generate the chart image. Charts generated without modification correspond to the correct versions. Operating at the annotation level also enables compatibility with other visualization libraries.
@@
Inappropriate axis range The axis range is either too broad or too narrow to accurately visualize the data, allowing changes to be minimized or maximized depending on the author’s intention. [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Inappropriate item order The items are arranged in an unconventional order, misleading the audience or creating confusion. [Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
Truncated axis The axis does not start from zero or is truncated in the middle, resulting in an exaggerated difference between the two bars. [Uncaptioned image][Uncaptioned image]
@@
Evaluation and filtering process.
We perform chart QA on each correct chart and its misleading counterpart generated by our rule-based system and evaluate performance using relaxed accuracy Masry et al. (2022); Methani et al. (2020). We use three instruction-tuned MLLMs selected based on ChartQA test-set performance (Masry et al., 2022): QwenVL 2.5-32B, InternVL 3.0-38B, and KimiVL-A3B. We retain instances where the majority of models answer correctly on the original chart but incorrectly on the misleading chart. A consistency filter ensures that errors are attributable to the misleader: numeric answers must have a standard deviation below 0.5, while textual answers must share a majority identical incorrect response. The final misleading answer is obtained by averaging incorrect numeric responses or taking the majority vote for textual responses.
Cross-domain extension.
We apply the same pipeline to two additional datasets to extend our corpus to new domains. First, we use ChartQA (Masry et al., 2022), which contains charts from real-world sources (Statista, Pew Research Center, Our World in Data, and the OECD). Due to inconsistent or missing annotations, we merge all instances into a single test set (Appendix C.2). Second, we use ChartX (Xia et al., 2025), which includes chart types that can be directly converted into structured data and covers different domains (commerce, industry, society, culture, and lifestyle).
@@ 5 Experiments
Figure 4: Average accuracy on AttackViz. Top: Results by model. Bottom: Results by misleader. Colors denote dataset and evaluation setting: PlotQA (Accuracy on correct charts, Accuracy on misleading charts), ChartQA (Accuracy on correct charts, Accuracy on misleading charts), and ChartX (Accuracy on correct charts, Accuracy on misleading charts).
@@ 5.1 Experimental setup
We perform all experiments on the test splits of AttackViz, derived from PlotQA, ChartQA, and ChartX. We evaluate both in-domain and cross-domain generalization. In the in-domain setting, demonstrations and test instances come from PlotQA. In the cross-domain setting, demonstrations are selected from the PlotQA train split, while test instances are drawn from ChartQA and ChartX. For each test instance, the Demonstration Selection module retrieves the most relevant training examples as demonstrations.
@@
Following prior work on MLLM vulnerabilities to misleading charts Tonglet et al. (2025a), we evaluate 16 open-weight instruction-tuned models: Ovis-2.5 (2B, 9B) Lu et al. (2025), InternVL-3.5 Wang et al. (2025) (1B, 2B, 4B, 8B, 14B, 38B), LLaVA-1.6 Liu et al. (2024) (7B, 13B, 34B), Qwen3-VL Yang et al. (2025) (2B, 4B, 8B, 32B), and LLaMA-4 Meta AI (2025) (17B-16E). Open-weight models are loaded using HuggingFace Transformers Wolf et al. (2019). We also evaluate three close models: GPT-4o Alexander et al. (2024), Gemini-2.5-Flash Comanici et al. (2025), and Claude-4.6-Sonnet Anthropic (2026). We perform inference on these close models using the OpenRouter API. [^2] https://openrouter.ai/ We exclude chart-specialized MLLMs, as recent general-purpose MLLMs outperform them Nguyen et al. (2026); Wang et al. (2025).
@@
We evaluate each model under two settings: (i) with the correct chart and (ii) with the misleading chart generated by ChartAttack. Following prior work Methani et al. (2020); Masry et al. (2022), we report relaxed accuracy as the primary metric. We also introduce two deception-rate metrics. Deception rate (originally correct) measures the percentage of instances where a model answers correctly on the correct chart but outputs the misleading answer on the misleading chart. Deception rate (originally incorrect) measures the percentage of instances where an incorrect answer is replaced by the misleading answer, indicating whether misleading charts reinforce existing errors. Achieving high deception rates is challenging because these metrics require an exact match with the misleading answer, meaning the Misleader-generator must anticipate the model’s behavior.
@@ 5.2 MLLM-based evaluation results
Figure 5: Average deception rate (DR) on AttackViz. Top: Results by model. Bottom: Results by misleader. Colors denote dataset and evaluation setting: PlotQA (DR on correct, DR on incorrect), ChartQA (DR on correct, DR on incorrect), and ChartX (DR on correct, DR on incorrect). Bars are stacked to represent conditional deception rates.
@@
All models perform worse on misleading charts than on correct charts, with accuracy drops ranging from 4.4 to 26.6 pp (17.2 pp on average). The degradation increases with model capability. Lower-performing models such as LLaVA-1.6 variants (28–44% accuracy on correct charts) show small declines of 4–10 pp. Mid-range models, including Claude-4.6, GPT-4o, Gemini-2.5, Qwen3-VL, and smaller InternVL variants (52–77%), drop by 12–21 pp. Higher-performing models such as InternVL-3.5 (14B/38B) and Ovis-2.5 (2B/9B) achieve 80–86% accuracy but exhibit larger declines of 22–27 pp. Conditional deception rates show that most errors arise when correct answers shift to attacker-generated misleading answers (11.2% on average), whereas originally incorrect answers rarely change (1.7%). The impact varies across model families. LLaVA-1.6 drops increase modestly from 4.9 pp (7B) to 9.8 pp (34B) despite the large size difference. InternVL-3.5 shows stronger scaling effects, with drops rising from 15.8 pp (1B) to 21.3 pp (8B) and 26.6 pp (38B), though not strictly monotonically. For example, InternVL-3.5 14B (24.4 pp) drops more than InternVL-3.5 26B (22.8 pp). Cross-family comparisons reveal similar vulnerabilities across architectures, such as InternVL-3.5 14B (24.4 pp) and Ovis-2.5 9B (23.3 pp). Other strong multimodal models including GPT-4o, Gemini-2.5, Claude-4.6, and Qwen3-VL show substantial but intermediate drops.
At the misleader level, perceptual manipulations produce the strongest effects. Stacked charts, 3D charts, and inappropriate log scales reduce accuracy to 24.6%, 34.2%, and 42.1%, corresponding to drops of 41.5 pp, 30.6 pp, and 18.8 pp, and higher deception rates on originally correct answers (20.0%, 10.7%, and 6.9%). Misrepresentation and inverted axes cause moderate declines (18.8 pp and 19.4 pp) with deception rates of 9.6% and 15.6%. Inappropriate line charts have a smaller impact (9.0 pp drop, 6.5% deception), while ineffective color schemes have minimal effect (0.4 pp increase, 1.2% deception). Dual axes lead to only a 1.4 pp overall decline; however, when applied effectively, they can still produce misleading answers generated by ChartAttack (10.9% deception).
@@
Consistent with in-domain results, accuracy on misleading charts drops in ChartQA and ChartX by 4.2–19.1 pp across models, with average declines of 11.5 pp and 12.3 pp, respectively. High-performing models are not immune: InternVL-3.5 (14B/38B), Ovis-2.5 9B, GPT-4o, Gemini-2.5, and Claude-4.6 also show substantial degradation. Conditional deception rates remain relatively low compared with in-domain experiments, averaging 11.7% and 14.9% for originally correct answers and 2.7% and 1.9% for originally incorrect answers on ChartQA and ChartX.
At the technique level, several trends from the in-domain experiments persist. 3D remains the most impactful technique, reducing accuracy to 27.9% and 22.7% (drops of 55.1 pp and 61.4 pp) on ChartQA and ChartX, with deception rates on originally correct answers of 4.2% and 5.0%. Misrepresentation follows, with 56.9% and 54.1% accuracy (-20.6 pp and -23.0 pp) and deception rates of 6.1% and 6.3%. Inappropriate stacked bars also remain effective, yielding 59.3% and 55.9% accuracy (-17.2 pp and -20.5 pp) and deception rates of 3.7% and 4.3%. In contrast, Dual axis and Ineffective color scheme remain largely ineffective, producing negligible changes (+0.2 pp and 0.0 pp) and small drops (1.8 pp and 2.6 pp). Inappropriate line charts, log scales, axis ranges, and truncated axes show a different pattern: although effective in PlotQA, they cause only modest drops in ChartQA and ChartX, suggesting their impact depends more strongly on dataset characteristics. These results indicate that some misleaders generalize across domains, while others are more sensitive to chart and question semantics. We analyze performance drops across families, chart types, and misleaders in both settings in Appendix E.